OpenStudy (anonymous):

4√2 + 7√2 - 3√2 Can you show me how to do this? Home schooled and don't understand this lesson.

OpenStudy (across):

Factor

OpenStudy (anonymous):

sqrt(2)(4+7-3)= 8(SQRT(2))

OpenStudy (wasiqss):

as the value in under root is same so take them common 8underroot 2 is answer

OpenStudy (anonymous):

Just add the numbers in front of the square root since they all have the same number inside the radical.. so it would be \[8\sqrt{2}\]

OpenStudy (hoblos):

=√2(4+7-3) = 8√2

OpenStudy (anonymous):

Thank you all, I understand it a lot better thanks to Tyler1992 for the good explanation. (:

OpenStudy (anonymous):

That only works if they have the same number under the radical though.. But you're welcome.